the ordered pair (8,6) is a solution for which equations ?
y=6x
y=8x
y=(2/3)x
y=(3/4)x

Answer by Edwin McCravy(20055)

(8,6) means to let x=8 and y=6
----------------------------------
y = 6x   Substitute x=8 and y=6
6 = 6(8)
6 = 48
That's false because 6 does not equal 48.  So (8,6) is not a solution to y=6x.
----------------------------------
y = 8x   Substitute x=8 and y=6
6 = 8(8)
6 = 64
That's false because 6 does not equal 64.  So (8,6) is not a solution to y=8x.
either.
----------------------------------
y = (2/3)x   Substitute x=8 and y=6
6 = (2/3)(8)
6 = 16/3
That's false because 6 does not equal 16/3.  So (8,6) is not a solution to y=(2/3)x.
----------------------------------
y = (3/4)x   Substitute x=8 and y=6
6 = (3/4)(8)
6 = 6
That's true because 6 does indeed equal 6.  So (8,6) is indeed a solution to y=(3/4)x. 
So this is the correct answer. 

Edwin